What is the pH of the solution if o.017 mol of Hcl is added to a buffer containing 0.015 mol of RcooH and 0.039 mol of RCOONa (source of RCoo Ka 2.1 E-5 Answer: Check Oxalic acid HOOCCOOH, is a diprotic acid with K 0.056 and K 1.5x10 Determine the equilibrium [HOOCCOOH] in a solution with an initial concentration of 0.15 M oxalic acid. HOOCCOOH(aq HOOCCOO (aq) H (aq) HOOCCOO (aq) OOCCOO (aq) H (aq) Answer
Answer #1

(a)

Given:

Ka = 2.1*10-5, so pKa = -log(Ka) = 4.67

Using the following equation:

pH = pKa + log(moles of salt/moles of acid)

After addition of HCl, net moles of acid increase and moles of salt decrease.

So, at equilibrium:

moles of acid = 0.015 + 0.017 = 0.032

moles of salt = 0.039 - 0.017 = 0.022

Putting values in the above equation, we get:

pH = 4.67 + log(0.022/0.032) = 4.5

(b)

For the diprotic acid, the second dissociation constant is almost negligible, so the [HOOCCOOH] dissociates into the solution almost completely through the first reaction.

Consider the reaction below:

H2C2O4 ----> H+ + HC2O4-

Initial 0.15 0 0

Eqb 0.15-x x x

Ka1 = ([H+][HC2O4-]) / [H2C2O4] = x2/(0.15-x) = 0.056

Solving above equation, we get:

x = 0.0678

So, at eqb.,

[HOOCCOOH] = 0.15 - 0.0678 = 0.0822 M